OpenStudy (anonymous):

i am wonder if i did this correct. (6x+7)(3x+2) = 13?

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

i got 13

OpenStudy (anonymous):

18X^2+33X+1=0

OpenStudy (anonymous):

Then no. You have x terms in there that aren't gonna go away.

OpenStudy (anonymous):

Are you given (6x+7)(3x+2) and asked to simplify? Or is it an equation?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

Quadratic formula, anyone?

OpenStudy (anonymous):

Right. So you have to distribute.. \[(6x+7)(3x+2) = 6x(3x+2) + 7(3x+2)\] And then distribute again.

OpenStudy (anonymous):

@dheeraj how did u get 0?

OpenStudy (anonymous):

It's not an equation. You can't solve it using the formula.

OpenStudy (anonymous):

Woops, my bad wasn't really paying attention.

OpenStudy (anonymous):

dheera thought you were working with the equation (6x+7)(3x+2) = 13

OpenStudy (anonymous):

But you aren't, so his solution is not what you want.

OpenStudy (anonymous):

i guess i right after all

OpenStudy (anonymous):

No you aren't

OpenStudy (anonymous):

i wasnt ....so the answer is not 13 ?

OpenStudy (anonymous):

No. I already said that, and explained how to find the expanded form.

OpenStudy (anonymous):

YEP I THOUGHT OF QUADRATIC EQUATION

OpenStudy (radar):

If you review this thread it was not an equation. It was simply a multiplication problem. So you won't have to use quadratic formula.

OpenStudy (anonymous):

exactly let me give u an example: (5x+2) (7x-3) = 35x^2-x-6

OpenStudy (radar):

Yes and from what polpak provided you will get;\[18x ^{2}+33x+14\] provided i multiplied correctly

OpenStudy (radar):

Did you get that also?

OpenStudy (anonymous):

yes but i am trying to figure out how is that possiable

OpenStudy (radar):

Some people use what they call FOIL. I like to do it the old fashion way 6x+7 times 3x+2 ---------------- 12x+14 18x^2+ 21x ----------------- 18x^2+33x+14

OpenStudy (anonymous):

i got it did i guess i didnt mine the wrong way

OpenStudy (anonymous):

I don't like foil or the 'old fashioned' way. I just distribute. \[(6x+7)(3x+2) \]\[= 6x(3x+2) + 7(3x+2)\]\[=18x^2 + 12x + 21x + 14\]\[=18x^2 + 33x + 14\]

OpenStudy (radar):

That looks like a simpler way than FOIL

OpenStudy (anonymous):

And it works when there's more than 2 terms!

OpenStudy (anonymous):

thanks guys

OpenStudy (radar):

Very good, a simple and to me a novel method. I never did like FOIL

OpenStudy (anonymous):

It's not really novel.. You do it all the time for monomials. It's just a continuation of standard distribution. \[3b(2c+d) = 3b(2c) + 3b(d)\] But if b = (f+g) We still get the right thing

OpenStudy (anonymous):

ok sounds good

OpenStudy (anonymous):

ok i got another (4x+1)(2x^2+x+1)